OpenStudy (anonymous):

A businessman bought a dealership that is incurring a loss of $500,000 a year. He decided to strategize in order to turn the business around. In addition to the $500,000 annual loss, his fixed cost for running the dealership on a monthly basis is $5,000. The number of cars sold per week and their probabilities mimic the outcomes of three coins being flipped. The number of cars sold in a week was observed to be the same as the number of tails that appear in a three coin flip. See the distribution: Number of Tails 0 |1 |2 |3 probablity 1/8 | 3/8 | 3/8 | 1/8

OpenStudy (anonymous):

Given that there are 52 weeks in a year, what is the expected revenue per car (rounded to the nearest dollar) that has to be made in order to break even in the first year?

OpenStudy (anonymous):

A. $4,308 B. $7,179 C. $5,385 D. $10,769 E. $3,590

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

5000*12 = 60 000 for the year, add in the loss we get 560 000 that needs to be covered in a year; divided by 52 gives a weekly breakeven amount of say: 10 770 ----------------------- the expected probability i believe is a weighted average.\[0(1/8)+1(3/8)+2(3/8)+3(1/8)=12/8=1.5\] so he should expect to sell 1.5 cars on average per week. 10770/1.5 = 7180 per car

OpenStudy (amistre64):

but thats just an idea, i cant verify it tho :/